Find the inverse Laplace of {2s+5/s^2+25}?

A) 2Sin5t+Cos5t

B) Cos5t-2Sin5t

C) 2Cos5t+Sin5t D) 2Cos5t-Sin5t

1
Expert's answer
2022-02-14T17:13:54-0500
"L^{-1}(\dfrac{2s+5}{s^2+25})=2L^{-1}(\dfrac{s}{s^2+25})+L^{-1}(\dfrac{5}{s^2+25})"

"=2\cos(5t)+\sin (5t)"

C) 2Cos5t+Sin5t